OpenStudy (anonymous):

Standard Rental Plan, they can rent a car for $52/day plus 23 cents/mile Deluxe Rental Plan, they can rent a car for $80/day with unlimited mileage. Write an equation that represents the Standard Rental Plan.

OpenStudy (jazzyfa30):

52d+23m=x

OpenStudy (jack1):

if x is the amount paid for the length of the rental and d = the number of days the car is rented for and m = the number of miles travelled in the car then Standard Rental Plan - 52d + 0.23m = x or Deluxe Rental Plan - 80d = x

OpenStudy (jazzyfa30):

the reason being is because it is $52 a day so because we don't know how long they are going to have it we writ it as 52d then because it is $.23 per mileage we don't know how far they are going so we do 23m so when you go to fill in your variables your x will be the amount of money they will spend so that is ho w you get the equation of 52d+23m=x